OpenStudy (anonymous):

I don't understand what I am supposed to do I think I am supposed to turn this into slope intercept form but, I don't know how. Will someone please help me? What is the equation that represents a line passing through the point 2, -4 with a slope of 1/2?

OpenStudy (studygurl14):

slope-intercept form: y=mx + b where m = slope, and b = y-intercept

OpenStudy (studygurl14):

They already give you the slope, so what's the first step?

OpenStudy (anonymous):

I am not sure sorry I'm terrible at slope intercept form. :(

OpenStudy (anonymous):

y = mx + b slope(m) = 1/2 (2,-4)...x = 2 and y = -4 now we sub...and since we already have slope(m), we need to find y intercept(b) -4 = 1/2(2) + b -4 = 1 + b -4 - 1 = b -5 = b so your equation is : y = 1/2x - 5

OpenStudy (studygurl14):

quote: "slope of 1/2" If m= slope in the equation y=mx + b, what can you replace "m" with?

OpenStudy (anonymous):

1/2?

OpenStudy (studygurl14):

correct.

OpenStudy (anonymous):

if slope is m then what is b? and what is y? What do these represent?

OpenStudy (studygurl14):

b is the y-intercept. But you should look at @kelliegirl33 's post. She told you step by step what to do, and gave you the answer. :)

OpenStudy (anonymous):

ok I am just trying to do the same thing she did on paper so that I can understand how to do this I am still confused a bit not much though thanks

OpenStudy (anonymous):

I think I got it now thank you both! :) I am just a little confused at the end of this problem but, I will keep practicing thank you!

OpenStudy (anonymous):

good luck...with enough practice, you will soon be helping others learn this